# Topology-Aware Conformal Prediction for Stream Networks

### v1

@@ 5 Theoretical Analysis
    Optimal Efficiency: We establish that STACI maximizes predictive efficiency by reducing the uncertainty set volume, justifying the need for accurate covariance estimation in spatio-temporal stream networks (Theorem 5.5).
@@
    Validity Guarantees under Stationarity and Adaptation to Distribution Shifts: We prove that STACI ensures valid conditional coverage under stationary assumptions (Theorem 5.4) and extend the framework to handle non-stationary settings via an ACI adjustment, ensuring approximate average coverage (Proposition 5.6).
Our analysis is based on the Mahalanobis distance framework in Equation 4.1, which enables the construction of arbitrary ellipsoidal uncertainty sets, providing greater flexibility in evaluating various nonconformity scores. For example, standard CP with spherical uncertainty sets arises as a special case when $A$ is an identity matrix. Another instance is the approach in Xu and Xie [39], where $A$ is set as the sample covariance matrix.
@@ 5.1 Coverage Validity
We analyze the conditional coverage validity of the proposed method. Consider the additive error model described in Equation 3.1 where the errors, $\epsilon_{t}$, are i.i.d.. We introduce the following assumption and, for simplicity, denote the nonconformity score $\epsilon_{t}^{\top}A\epsilon_{t}$ as $s_{t}$.
Assumption 5.1 (Estimation quality).
There exists a sequence $\{\nu_{n}\}$, $n\geq 1$ such that $\frac{1}{n}\sum_{t=T-n+1}^{T}||\epsilon_{t}-\hat{\epsilon}_{t}||^{2}\leq\nu^{2}_{n},||\epsilon_{T+1}-\hat{\epsilon}_{T+1}||\leq\nu_{n}$
@@
The assumption ensures that the prediction error is bounded by $\nu_{n}^{2}$. In many practical estimator, the $\nu_{n}$ vanishes as $n\to\infty$, indicating improved estimation accuracy with larger sample sizes.
Assumption 5.2 (Convergence of $A_{n}$).
The sequence $\{A_{n}\}$ associated with the nonconformity score is assumed to converge to a fixed matrix $A$ as $n$ increases, with an upper-bounded convergence rate $o(g(n))$,
$\|A_{n}-A\|\;=\;o\bigl{(}g(n)\bigr{)}.$
Additionally, there exists a constant $r>0$ such that $\|A_{n}\|>0$ and $\|A\|\leq r$.
@@
When designing nonconformity scores, the matrix $A$ can be chosen to either remain constant or converge to a fixed matrix. The flexibility in selecting $A$ allows for adaptability across different application scenarios. For example, if the true covariance matrix of the error $\epsilon$ is known, $A$ can be set as its inverse. Alternatively, if only sample estimates are available, $A$ can be chosen as the inverse of the estimated sample covariance matrix of $\epsilon$, provided it converges under appropriate tail behavior conditions [34]. The major difference between between different choices of $A_{n}$ lies in their respective convergence rates.
Assumption 5.3 (Regularity conditions for $s_{t}$ and $\epsilon_{t}$).
Assume that the cumulative distribution function (CDF) of the true nonconformity score, $F_{s}(x)$, is Lipschitz continuous with a constant $L>0$. Suppose there exist constants $\kappa_{1},\kappa_{2}>0$ such that:
$\|\epsilon_{t}\|\leq\kappa_{1}I~{}\text{almost surely, and}~{}\mathrm{Var}[\|\epsilon_{t}\|^{2}]\leq\kappa_{2}I.$
Theorem 5.4 (Validity).
@@
$\left|\mathbb{P}(Y_{T+1}\in\mathcal{C}_{T+1}(\alpha)|X_{T+1}=x)-(1-\alpha)\right|$
$\leq$ $(4L+2L\sqrt{\omega}+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+\frac{\log(16n)}{n},$
@@
The finite-sample bound on the coverage gap is directly influenced by the estimation quality and the convergence rate of $A_{n}$, which is given by $\max(\mathcal{O}(\frac{\log n}{n}),\mathcal{O}(\nu_{n}),\mathcal{O}(\sqrt{g(n)}))$. In general, reducing the coverage gap requires high-accuracy estimations (i.e., a rapidly vanishing $\nu_{n}$) and a well-chosen nonconformity score matrix $A_{n}$ that converges quickly.
Theorem 5.4 highlights the importance of incorporating topology-based estimators in STACI. Relying solely on the sample covariance matrix often leads to coverage gaps in finite samples, undermining validity. In contrast, the topology-based matrix acts as a covariance estimator with topology-informed regularization, generally achieving faster convergence than the sample covariance estimator. A hybrid approach that combines both estimators provides an optimal trade-off between validity and efficiency.
@@ 5.2 Prediction Efficiency
We now analyze the efficiency of STACI. The predictive efficiency is evaluated based on the volume of the prediction set in $I$-dimensional space, defined as
$V(A,r)=\frac{\pi^{I/2}}{\Gamma\left(\frac{I}{2}+1\right)}\cdot r^{I/2}\cdot\det(A)^{-1/2}.$ (5.1)
The radius of the prediction set is determined by the $(1-\alpha)$-th quantile of the empirical CDF, computed from $n$ data points in the calibration dataset. This radius is denoted as $Q_{1-\alpha}(\{\hat{\epsilon}_{t}^{\top}A\hat{\epsilon}_{t},t\in\mathcal{D}_{\text{cal}}\})$ In the ideal case where $\hat{f}(X_{t})=f(X_{t})$ and $\hat{\epsilon}_{t}=\epsilon_{t}$, minimizing inefficiency reduces to solving the following optimization problem:
$\min_{A\succ 0}V(A,\hat{Q}_{1-\alpha}(\{\epsilon_{t}^{\top}A\epsilon_{t},t\in[n]\})).$ (5.2)
Since computing the quantile of the empirical CDF directly can be complex, we approximate the solution in Equation 5.2 by replacing the empirical CDF with the true CDF. This approximation is justified by the Glivenko-Cantelli Theorem [31], which ensures that
$\lim_{n\to\infty}\hat{Q}_{1-\alpha}(\{\epsilon_{t}^{\top}A\epsilon_{t},t\in[n]\})=Q_{1-\alpha}(\epsilon^{\top}A\epsilon),$
where $Q_{1-\alpha}(\cdot)$ denotes the $1-\alpha$ quantile of $\epsilon^{\top}A\epsilon$ and is assumed to be continuous.
In the limiting case, we formulate the following minimization problem, presented in Theorem 5.5, and use its solution as the guiding criterion for selecting the matrix $A$:
Theorem 5.5 (Efficiency).
The optimal solution to the minimization problem is given by:
$A_{*}\coloneqq\arg\min_{A\succ 0}V(A,Q_{1-\alpha}(\epsilon^{\top}A\epsilon)),$ (5.3)
@@
Since the optimization problem is invariant to scalar rescaling (i.e., $A$ and any positive scalar multiple $cA$, where $c>0$, yield the same mathematical solution), the primary focus is on identifying the structural form of $A$. For computational tractability, additional constraints, such as bounding the matrix norm $\|A\|\leq 1$, can be imposed without loss of generality. While the assumption that $\epsilon$ follows a Gaussian distribution simplifies analysis, the result can be extended to broader distributions that satisfy appropriate tail-bound conditions.
Theorem 5.5 underscores the importance of selecting $A$ optimally in Equation 4.1 and highlights that accurately estimating the inverse of the error covariance matrix is key to minimizing inefficiency in CP. In practice, designing an optimal $A_{*}$ is often challenging due to empirical limitations. For example, the estimated residuals $\hat{\epsilon}_{t}$ may deviate significantly from the true errors $\epsilon_{t}$. Additionally, when the sample size $n$ is small, the empirical CDF may differ considerably from the true CDF. Despite these challenges, constructing $A$ based on an estimate of the inverse covariance matrix offers substantial improvements in high-dimensional settings compared to CP methods that ignore variable dependencies, such as those that set $A$ as the identity matrix.
5.3 Adaptively Adjusting the Confidence Level
We present the analysis of the average coverage guarantee of STACI without any assumption about $\epsilon_{t}$. The proof follows from Proposition 4.1 in [10].
Proposition 5.6.
$\lim_{n^{\prime}\to\infty}\sum_{t\in\mathcal{D}_{\text{test}}}\mathds{1}\{Y_{t}\notin\mathcal{C}(X_{t};\alpha_{t})\}\big{/}n^{\prime}=\alpha.$
While Proposition 5.6 provides a weaker coverage guarantee compared to Theorem 5.4, it offers broader applicability, remaining valid even in adversarial online settings. Empirical results suggest that when the error process exhibits minimal distribution shift and the assumptions of Theorems 5.5 and 5.4 are only slightly violated, STACI maintains the predefined coverage level ($\gamma=0.01$) while achieving efficient prediction sets. However, when $\gamma>0$, Proposition 5.6 does not ensure a finite-sample coverage gap. Understanding this limitation and developing methods to control the finite-sample coverage gap presents an interesting direction for future research.
@@ 6 Experiments
To demonstrate the suitability of our proposed method, STACI, for stream networks, we evaluate its performance on both synthetic data with a stationary covariance matrix and real-world data with time-varying covariance. By default, the first $60$% of observations are used for training, the calibration set consists of the most recent $n=300$ observations, and the test contains the sequentially revealed observations $n^{\prime}=5000$ in simulation and $n^{\prime}=3000$ in real study. The weighting factor $\lambda$ is set to $0.5$. The desired confidence rate $\alpha$ is fixed at $0.95$. Our method is compared against five baselines: ($i$) Sphere: Spherical confidence set, where the covariance matrix is an identity matrix. In another word, the prediction error at different locations are not considered to have correlations. ($ii$) Sphere-ACI ($\gamma=0.01$): Spherical confidence set with adaptive conformal inference (ACI). ($iii$) Square: Square confidence set. This equals to computing different nonconformity scores for each dimension, and then calibrate accordingly. ($iv$) GT: Ellipsoidal confidence set using the ground-truth covariance matrix. ($v$) MultiDimSPCI: Ellipsoidal confidence set using the sample covariance matrix [40]. We consider both validity and efficiency to evaluate the uncertainty quantification performance: ($i$) Coverage quantifies the likelihood that the prediction set includes the true target, i.e.,
$\text{Coverage}:=\sum_{t\in\mathcal{D}_{\text{test}}}\mathds{1}\{Y_{t}\notin\mathcal{C}(X_{t};\alpha_{t})\}\big{/}n^{\prime}.$
($ii$) Efficiency is evaluated based on the size (or volume) of the prediction set, with smaller sets indicating higher efficiency. The volume of the prediction set, $\text{Vol}(\mathcal{C}(X_{t};\alpha_{t}))$, is measured by the size of the ellipsoid determined by $A$, as specified in Equation 5.1. Formally,
$\text{Efficiency}:=\sum_{t\in\mathcal{D}_{\text{test}}}\Big{(}\text{Vol}(\mathcal{C}(X_{t};\alpha_{t}))\Big{)}^{1/I}\big{/}n^{\prime}.$
An optimal method should achieve the predefined coverage with high efficiency.
Figure 2: Comparison of methods on synthetic datasets with different tail-up parameters $\Theta$ over coverage ($x$-axis) and efficiency ($y$-axis). Each method is evaluated over $10$ different seeds, with the circle size representing variance of coverage. The pre-determined coverage threshold of $95$% is indicated by a gray dotted line, where methods to the right achieve the desired coverage. Beyond meeting this threshold, better methods should also be positioned closer to the upper-right corner, as indicated by the red arrow, reflecting both higher coverage and greater efficiency.
@@ 6.1 Simulation
In this section, we conduct simulation experiments on synthetic data generated by a tail-up model. Specifically, we follow [24] and construct the stream network as shown in Figure 1. The details of synthetic network is provided in Appendix C.1. We generate the observation of site $u$ at time point $t$ by simulating stochastic integration from all upstream points $r\in\wedge u$ to downstream point $u$ according to Equation 4.3, where we set $\mu_{t}(u)=\sum_{i=1}^{w}\theta_{i}Y_{t-i}(u)$ following the $\text{AR}(w)$ structure and $m(\Delta)=\exp(-\Delta)$ as the exponential moving average function. The process is repeated until $5000$ time steps. This experiment simulates the stream network data without any misspecification.
@@
In synthetic data, the prediction model $f$ is simply a linear regression model. We first estimate parameters of in $\text{AR}(w)$ structure, i.e., $\Theta=(\theta_{i})_{i\in[w]}$, through linear regression and then parameters in Equation 4.4, $\phi$ and $\sigma^{2}$ through $\ell_{1}$-loss. Parameters of $\Theta=(0,0)$ and $\Theta=(0.7,0.3)$ are selected for data generation. When $\Theta=(0,0)$, the observations consist of pure noise, thus stationary; when $\Theta=(0.7,0.3)$, the process resembles a second-order autoregressive model.

### v2

@@ 5 Theoretical Analysis
    Optimal Efficiency: We establish that STACI maximizes predictive efficiency by reducing the uncertainty set volume, justifying the need for accurate covariance estimation in spatio-temporal stream networks (Theorem 2).
@@
    Validity Guarantees under Stationarity and Adaptation to Distribution Shifts: We prove that STACI ensures valid conditional coverage under stationary assumptions (Theorem 1) and extend the framework to handle non-stationary settings via an ACI adjustment, ensuring approximate average coverage (Proposition 1 in Section C of the Appendix).
Our analysis is based on the Mahalanobis distance framework in eq.˜2, which enables the construction of arbitrary ellipsoidal uncertainty sets, with greater flexibility for nonconformity scores. For example, standard CP with spherical uncertainty sets arises as a special case when $A$ is an identity matrix. Another instance is the approach in xu2023conformal, where $A$ is set as the sample covariance matrix.
@@ 5.1 Coverage Validity
We analyze the conditional coverage validity of the proposed method. Consider the additive error model described in eq.˜1 where the errors, $\epsilon_{t}$, are i.i.d.. We introduce the following assumption and, for simplicity, denote the nonconformity score $\epsilon_{t}^{\top}A\epsilon_{t}$ as $s_{t}$.
Assumption 1 (Estimation quality).
[…]
@@
The assumption ensures that the prediction error is bounded by $\nu_{n}^{2}$. For many estimators, the $\nu_{n}$ vanishes as $n\to\infty$, indicating improved estimation accuracy with larger sample sizes chen1999improved.
Assumption 2 (Convergence of $A_{n}$).
The sequence $\{A_{n}\}$ associated with the nonconformity score converges to a fixed matrix $A$ as $n$ increases, with an upper-bounded convergence rate $o(g(n))$, i.e. $\|A_{n}-A\|\;=\;o\bigl(g(n)\bigr).$ Additionally, there exists a constant $r>0$ such that $\|A\|\leq r$.
@@
When designing nonconformity scores, the matrix $A$ can be chosen to either remain constant or converge to a fixed matrix. The flexibility in selecting $A$ allows for adaptability across different scenarios. For example, if the true covariance matrix of the error $\epsilon$ is known, $A$ can be set as its inverse. Alternatively, if only sample estimates are available, $A$ can be chosen as the inverse of the estimated sample covariance matrix of $\epsilon$, provided it converges under proper tail behavior conditions vershynin2012close. The major difference between choices of $A_{n}$ lies in the respective convergence rates.
Assumption 3 (Regularity conditions for $s_{t}$ and $\epsilon_{t}$).
Assume that the cumulative distribution function (CDF) of the true nonconformity score, $F_{s}(x)$, is Lipschitz continuous with a constant $L>0$. Suppose there exist constants $\kappa_{1},\kappa_{2}>0$ such that $\|\epsilon_{t}\|\leq\kappa_{1}I~\text{almost surely, and}~\mathrm{Var}[\|\epsilon_{t}\|^{2}]\leq\kappa_{2}I.$.
Theorem 1 (Validity).
@@
$\left|\mathbb{P}(Y_{T+1}\in\mathcal{C}_{T+1}(\alpha)|X_{T+1}=x)-(1-\alpha)\right|\leq(4L+2L\sqrt{\omega}+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+\frac{\log(16n)}{n},$
@@
[…]
Theorem 1 highlights the importance of incorporating topology-based estimators in STACI. Relying solely on the sample covariance matrix often leads to coverage gaps in finite samples, undermining validity. In contrast, the topology-based matrix acts as a covariance estimator with topology-informed regularization, generally achieving faster convergence than the sample covariance estimator. A hybrid approach that combines both estimators provides an optimal trade-off between validity and efficiency.
@@ 5.2 Prediction Efficiency
We evaluate the efficiency of STACI via the volume of the prediction set in $I$-dimensional space, defined as $V(A,r)=\frac{\pi^{I/2}}{\Gamma\left(\frac{I}{2}+1\right)}\cdot r^{I/2}\cdot\det(A)^{-1/2}.$ The radius of the prediction set is determined by the $(1-\alpha)$-th quantile of the empirical CDF, computed from $n$ data points in the calibration dataset. This radius is denoted as $\hat{Q}_{1-\alpha}(\{\hat{\epsilon}_{t}^{\top}A\hat{\epsilon}_{t},t\in\mathcal{D}_{\text{cal}}\})$. In the ideal case where $\hat{f}(X_{t})=f(X_{t})$ and $\hat{\epsilon}_{t}=\epsilon_{t}$, minimizing inefficiency reduces to: $\min_{A\succ 0}V(A,\hat{Q}_{1-\alpha}(\{\epsilon_{t}^{\top}A\epsilon_{t},t\in[n]\})).$ To simplify computation, we approximate the empirical quantile with the true quantile, justified by the Glivenko-Cantelli Theorem tucker1959generalization, which ensures the convergence $\lim_{n\to\infty}\hat{Q}_{1-\alpha}(\{\epsilon_{t}^{\top}A\epsilon_{t},t\in[n]\})=Q_{1-\alpha}(\epsilon^{\top}A\epsilon),$ where $Q_{1-\alpha}(\cdot)$, the $1-\alpha$ quantile of $\epsilon^{\top}A\epsilon$ is assumed to be continuous.
In the limiting case, we formulate the following minimization problem, presented in Theorem 2, and use its solution as the guiding criterion for selecting the matrix $A$:
Theorem 2 (Efficiency).
There exists $0<\alpha_{0}<1$, such that when $\alpha<\alpha_{0}$, the optimal solution to the minimization problem is given by:
$A_{*}\coloneqq\arg\min_{A\succ 0}V(A,Q_{1-\alpha}(\epsilon^{\top}A\epsilon)),$ (7)
@@
[…]
Theorem 2 underscores the importance of selecting $A$ optimally in eq.˜2 and highlights that accurately estimating the inverse of the error covariance matrix is key to minimizing inefficiency in CP. In practice, designing an optimal $A_{*}$ is often challenging due to empirical limitations. For example, the estimated residuals $\hat{\epsilon}_{t}$ may deviate significantly from the true errors $\epsilon_{t}$. Additionally, when the sample size $n$ is small, the empirical CDF may differ considerably from the true CDF. Despite these challenges, constructing $A$ based on an estimate of the inverse covariance matrix offers substantial improvements in high-dimensional settings compared to CP methods that ignore variable dependencies, such as those that set $A$ as the identity matrix.
@@ 6 Experiments
To demonstrate the suitability of STACI, we evaluate its performance on both synthetic data with a stationary covariance matrix and real-world data with time-varying covariance. By default, the first 6060% of observations are used for training, the calibration set consists of the most recent $n=500$ observations, and the test contains the sequentially revealed observations $n^{\prime}=5000$ in simulation and $n^{\prime}=5000$ in real study. The desired confidence rate $\alpha$ is fixed at $0.95$. Our method is compared against the following conformal prediction and learning-based uncertainty quantification baselines: ($i$) Sphere: Spherical confidence set, where the covariance matrix is an identity matrix. In other words, the prediction error at different locations are not considered to have correlations. ($ii$) Sphere-ACI ($\gamma=0.01$): Spherical confidence set with adaptive conformal inference (ACI). (iiiiii) Square: Square confidence set. This equals to computing different nonconformity scores for each dimension, and then calibrate accordingly. ($iv$) GT: Ellipsoidal confidence set using the ground-truth covariance matrix. ($v$) MultiDimSPCI: Ellipsoidal confidence set using the sample covariance matrix (xuconformal), alongside its localized variant using the most recent observations, MultiDimSPCI (local). ($vi$) CopulaCPTS: Prediction region based on modeling the joint distribution of forecast errors with a copula function (sun2023copula). (viivii) DeepSTUQ: A Bayesian deep learning model that quantifies uncertainty in spatio-temporal graphs by using graph convolutions and Monte Carlo dropout (qian2023uncertainty).
[…]
(a) Comparison of STACI with baselines.
(b) Impact of $\lambda$ on STACI.
Figure 2: Experiment results of synthetic data: (a) Comparison of methods on synthetic datasets with different tail-up parameters $\Theta$ over coverage ($x$-axis) and efficiency ($y$-axis). (b) Trade-off between coverage and efficiency on synthetic data, where the higher the better performance.
@@ 6.1 Simulation
In this section, we conduct simulation experiments on synthetic data generated by a tail-up model. Specifically, we follow (peterson2010mixed) and construct the stream network as shown in Figure 1. The details of synthetic network is provided in Section D.1 of the Appendix. We generate the observation of site $u$ at time point $t$ by simulating stochastic integration from all upstream points $r\in\wedge u$ to downstream point $u$ according to eq.˜4, where we set $\mu_{t}(u)=\sum_{i=1}^{w}\theta_{i}Y_{t-i}(u)$ following the $\text{AR}(w)$ structure and $m(\Delta)=\exp(-\Delta)$ as the exponential moving average function. The process is repeated until $5000$ time steps. This experiment simulates the stream network data without any misspecification.
@@
In synthetic data, the prediction model $f$ is simply a linear regression model. We first estimate parameters of in $\text{AR}(w)$ structure, i.e., $\Theta=(\theta_{i})_{i\in[w]}$, through linear regression and then parameters in eq.˜29, $\phi$ and $\sigma^{2}$ through $\ell_{1}$-loss. Parameters of $\Theta=(0,0)$ and $\Theta=(0.7,0.3)$ are selected for data generation. When $\Theta=(0,0)$, the observations consist of pure noise, thus stationary; when $\Theta=(0.7,0.3)$, the process resembles a second-order autoregressive model. The weighting factor $\lambda$ is set to $0.6$.
Results
Our numerical results demonstrate that our method enhance the predictive efficiency significantly without sacrificing the coverage guarantee, by considering both sample-based and topology-based covariance. From Figure 2(a), we observe that CP methods employing ellipsoidal uncertainty sets tend to cluster towards the upper region of the plot, indicating higher efficiency compared to CP methods based on spherical or square uncertainty sets. Although MultiDimSPCI achieves the lowest efficiency, its coverage drops significantly below the required threshold, highlighting its instability when relying solely on the sample covariance matrix. This issue persists even in simulated data designed to align with its error process assumptions. Its local variant, designed to capture temporal correlations, provides only a marginal improvement and similarly fails to achieve the required coverage. In contrast, with $\lambda$ fixed at $0.6$, our method STACI is positioned near the upper-right corner alongside GT, which leverages the ground-truth covariance matrix. This suggests that our method achieves performance comparable to GT, balancing low efficiency (smaller volume) while maintaining the necessary coverage guarantees. Among all methods that surpass the coverage threshold, our method, STACI($\gamma=0.01$), demonstrates the best efficiency with the smallest variance, further reinforcing its robustness and effectiveness.
Figure 2(b) reveals a clear trade-off trend between coverage and efficiency: the higher $\lambda$, the confidence level rises, but efficiency is worse. This suggests that $\lambda$ should be carefully chosen: if too large, our method over-relies on topology and fails to adapt to covariance shift; if too small, it depends more on sample covariance matrices, which are purely data-driven and thus unstable, leading to a coverage drop. Nonetheless, no matter whether adapting confidence level, setting a larger $\lambda$ in STACI can efficiently increase coverage and maintain it near the pre-determined level, while only slightly increasing volume, which remains comparable to GT.
6.2 Real Data Study
We further conduct experiments on a real-world traffic dataset, Performance Measurement System (PeMS) chen2001freeway, which contains the data collected from the California highway network, providing $5$-minute interval traffic flow counts by multiple sensors, along with flow directions and distances between sensors. To model it into stream network, we also rely on CaDot_PEMS to check accurate road connection information. We select two highway forks, each equipped with $12$ sensors, named PeMS-G1 and PeMS-G2, and plot their locations and corresponding road segments in Figure 5 in the Appendix.
Experiment Configuration
We adopt Adaptive Graph Convolutional Recurrent Network (AGCRN) bai2020adaptive as the default backbone model $f$. To demonstrate STACI’s generality under post-hoc conformal prediction framework, we evaluate over alternative GNN backbones, including attention-based ASTGCN Guo_Lin_Feng_Song_Wan_2019 and continuous-time STGODE fang2021spatial. We set our default $\lambda=0.6$. For simplicity, we only use fixed weights with all equal values, without requiring any additional information. Multiple hyperparameter and ablation study are also provided over the key parameters in our framework: ($i$) $\lambda$ from 0 to $1$ with step of $0.02$; (iiii) $n=100,200,300,400,500$; (iiiiii) $\gamma=0$ or $0.01$.
Table 1: Comparison of Different CP Methods over Different GNN Models for two PeMS Datasets. Coverage below 94.5% is italicized. Inefficiency is reported as mean ± standard deviation, with the lowest value in bold.
Dataset Backbone Models AGCRN ASTGCN STGODE
CP Methods Coverage Efficiency $\downarrow$ Coverage Efficiency $\downarrow$ Coverage Efficiency $\downarrow$
PeMS-G1 Sphere $97.76$ $133.60\pm 12.79$ $96.64$ $128.23\pm 14.2$ $96.09$ $145.40\pm 7.79$
Sphere-ACI ($\gamma=0.01$) $95.26$ $108.93\pm 15.96$ $95.24$ $114.64\pm 19.54$ $95.03$ $136.36\pm 19.47$
Square $95.98$ $155.84\pm 23.92$ $96.41$ $155.62\pm 23.74$ $96.38$ $172.30\pm 23.51$
MultiDimSPCI 92.92 $73.82\pm 8.16$ 93.19 $74.68\pm 7.67$ 92.70 $88.48\pm 4.26$
MultiDimSPCI (local) $93.04$ $74.23\pm 8.27$ $93.57$ $75.18\pm 7.69$ $92.98$ $88.99\pm 4.26$
STACI ($\gamma=0$) $97.12$ $88.27\pm 21.73$ $96.76$ $88.1\pm 19.07$ $95.31$ $77.34\pm 7.09$
STACI ($\gamma=0.01$) $95.75$ 67.54 $\pm$ 10.94 $95.54$ 67.92 $\pm$ 10.22 $95.14$ 73.62 $\pm$ 9.83
PeMS-G2 Sphere $96.26$ $144.55\pm 14.22$ $95.64$ $130.69\pm 11.55$ $95.83$ $147.13\pm 14.68$
Sphere-ACI ($\gamma=0.01$) $95.03$ $133.63\pm 14.18$ $95.07$ $122.72\pm 16.67$ $95.14$ $122.34\pm 19.72$
Square $95.26$ $172.18\pm 19.43$ $95.37$ $160.60\pm 19.95$ $95.12$ $138.97\pm 17.78$
MultiDimSPCI 91.14 $101.42\pm 7.71$ 90.93 $92.12\pm 6.67$ 90.74 $107.23\pm 7.34$
MultiDimSPCI (local) $91.44$ $102.25\pm 7.90$ $91.12$ $92.84\pm 6.95$ $91.10$ $107.96\pm 7.58$
STACI($\gamma=0$) $95.39$ $73.36\pm 10.35$ $95.18$ $60.45\pm 9.22$ $95.33$ $77.14\pm 8.65$
STACI($\gamma=0.01$) $95.01$ 69.62 $\pm$ 12.85 $95.05$ 58.07 $\pm$ 9.83 $95.20$ 74.86 $\pm$ 9.63